The student dissolved 2.050 grams of acid in 50 mL of distilled water. He then titrated...

  1. The student dissolved 2.050 grams of acid in 50 mL of distilled water. He then titrated the unknown acid with a standardized NaOH solution. Unfortunately, the student over shot the end point and had to back titrate the solution with 2.17 mL of 0.98861 M HCl. Once the final end point was determined, the student had added a total of 29.15 mL of a 1.0324 M NaOH solution.

A) How many moles of OH- were used to reach the end point?

B) How many moles of H+ from HCl were used for back titration?

C) How many moles of H+ from the solid were neutralized?

D) What is the equivalent mass of the unknown acid?
Answer #1

A) No of mole of OH- used to reach the end point = moles of NaOH added - moles of HCl used to back titration.

   = 29.15*1.0324 - (2.17*0.98861)

   = 27.95 mmole

   = 0.02795 mole

B) No of mole of H+ were used to back titration = 2.17*0.98861

   = 2.1453 mmole

   = 0.0021453 mole

C) No of moles of H+ neutralised from solid = No of moleS of OH- added to reach endpoint

No of moles of H+ neutralised from solid = 0.02795 mole

D) Equivalent mass of the unknown acid = w/n

       = 2.050 / 0.02795

       = 73.35